Can you help me with this problem? Jill and Sarah are mowing a lawn together (with separate mowers). It will take them 2 1/2 hours to complete. If Jill (by herself) can mow the same lawn in 4 hours, how long would it take Sarah to mow the lawn by herself?
Answer by stanbon(75887)

Jill and Sarah are mowing a lawn together (with separate mowers). It will take them 2 1/2 hours to complete. If Jill (by herself) can mow the same lawn in 4 hours, how long would it take Sarah to mow the lawn by herself?
------
Together DATA: time = 5/2 hrs/job ; rate = 2/5 job/hr
-------
Jill DATA::::: time = 4 hrs/job ; rate = 1/4 job/hr
-------
Sarah DATA:::: time = x hrs/job ; rate = 1/x job/hr
=======================
Equation:
rate + rate = together rate
1/x + 1/4 = 2/5
-------
Multiply thru by 20x to get:
20 + 5x = 8x
3x = 20
x = 20/3 = 6 2/3 hrs (time for Sarah to do the job alone)
=============================================
Cheers,
Stan H.
==================